# Decomposition of Borel graphs and cohomology

### v2

@@ 3.1 Construction of trees
A family $\mathcal{C}$ of non-empty proper subsets of $X$ is a treeset on $X$ if
@@
    it is nested, i.e., for all $C,D\in\mathcal{C}$, either
    $C\cap D,\ C\cap\overline{D},\ \overline{C}\cap D\ \textup{or }\overline{C}\cap\overline{D}$
@@
    it is closed under complementation, i.e., if $C\in\mathcal{C}$, then $\overline{C}\in\mathcal{C}$ holds, and
@@
    it is finitely separating, i.e., for all $x,y\in X$, we have
    |$|\{C\in\mathcal{C}\mid x\in C,\ y\notin C\}|<\infty.$
The structure tree associated to a treeset is introduced by [Dun]. We will present a construction of it, which is based on [DD, Chapter II] but described from the viewpoint of ultrafilters in the sense of [Rol]. The proofs are given for the reader’s convenience.
@@
$C\subset D,\ C\subset\overline{D},\ \overline{C}\subset D\ \textup{or }\overline{C}\subset\overline{D}.$
@@
Next, we show that $(V_{\mathcal{C}},T_{\mathcal{C}})$ is acyclic. Let $\{u_{i}\}_{i=0}^{n}$ be a cycle of $T_{\mathcal{C}}$ with $n\geq 2$ such that the path $\{u_{i}\}_{i=0}^{n-1}$ is simple. Let $u_{0}\setminus u_{1}=\{C\}$. Then by Lemma 3.7, for $11\leq i\leq n-$ 1, we have $u_{i+1}\setminus u_{i}\neq\{C\}$ since $(u_{i+1},u_{i})\neq(u_{0},u_{1})$. This implies that $C\notin u_{i}$ for $1\leq i\leq n$ by the induction on $i$, which contradicts that $u_{n}=u_{0}$. Hence this graph is a tree. ∎
@@
3.2 Borel cutsets of Borel graphs
@@
Take a $G$-path $\{x_{i}\}_{i=0}^{n}$ from $x$ to $y$. If $C\in\mathcal{C}$ satisfies $x\in C$ and $y\notin C$, then we have $x_{i}\in\partial_{\textup{iv}}^{G}C$ for some $10\leq i\leq n-$ 1. By the previous lemma, there exist only finitely many $C\in\mathcal{C}$ satisfying this property. Hence the lemma is proved. ∎
The following is inspired by the proof of [CPTT, Proposition 3.3]:
@@
Take Borel subset $\mathcal{C}^{+}\subset\mathcal{C}$ so that |$\left|\left\{C,\overline{C}\right\}\cap\mathcal{C}^{+}\right|=1$ for every $C\in\mathcal{C}$. By [KST, Proposition 4.6], the Borel graph
[…]
@@
3.3 Structure trees for Borel graphs
@@
for every $10\leq i\leq n-$ 1. It suffices to verify that $x_{0}\neq x_{2n}$. Note that $\rho(x_{2i})\neq\rho(x_{2i+1})=\rho(x_{2i+2})$ for every $10\leq i\leq n-$ 1. We prove that
$d_{T_{\mathcal{C}}}(\rho(x_{0}),\rho(x_{2n}))=\sum_{i=0}^{n-1}d_{T_{\mathcal{C}}}(\rho(x_{2i}),\rho(x_{2i+2}))$ (3.2)
@@
4 Cohomology of graphs
Througout this section, let $R$ be a non-zero commutative ring and $(X,G)$ an abstract graph with uniformly bounded degrees.
4.1 Computaion of cohomology
@@
For $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩, set $\alpha_{\varphi}=1_{\textup{graph}\varphi}\in R_{G}$. Note that
@@
hold for $\varphi,\psi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ and $f\in l^{\infty}_{R}(X)$, where $\alpha_{\varphi*}f$ is the multiplication in the left $R_{G}$-module $l^{\infty}_{R}(X)$ as defined in Definition 1.10.
@@
If $G=\bigcup_{i=1}^{n}\textup{graph}\varphi_{i}$ with $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩, then $R_{G}$ is generated by the set $\{\alpha_{\varphi_{i}}\}_{i=1}^{n}\cup l^{\infty}_{R}(X)$ as a ring. Indeed, for every $k\in\mathbb{Z}_{\geq 0}$, there exists $\{\psi_{j}\}_{j=1}^{m}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ such that $G^{k}=\bigcup_{j=1}^{m}\textup{graph}\psi_{j}$ and each $\psi_{j}$ is obtained by composing elements of $\{\varphi_{i}\}_{i=1}^{n}$ at most $k$ times. Then for every $a\in R_{G}^{k}$, there exists $\{f_{j}\}_{j=1}^{m}\subset l^{\infty}_{R}(X)$ such that
@@
The algebra $R_{G}$ is projective as a left $l^{\infty}_{R}(X)$-module. Indeed, by Remark 4.2, we can take $\{\varphi_{i}\}_{i=1}^{\infty}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that for every $k\in\mathbb{Z}_{\geq 0}$, there exists $n\in\mathbb{Z}_{\geq 1}$ such that $G^{k}=\bigsqcup_{i=1}^{n}\textup{graph}\varphi_{i}$. Then we have $R_{G}=\bigoplus_{i=1}^{\infty}l^{\infty}_{R}(X)\cdot\alpha_{\varphi_{i}}$, where each $l^{\infty}_{R}(X)\cdot\alpha_{\varphi_{i}}$ is projective since we have
@@
$\cdots\to l^{\infty}_{R}(X)\otimes_{l^{\infty}_{\mathbb{Z}}(X)}P_{n}\to\cdots\to l^{\infty}_{R}(X)\otimes_{l^{\infty}_{\mathbb{Z}}(X)}P_{0}\to l^{\infty}_{R}(X)\to 0$ (4.1)
@@
for $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ and $f\in l^{\infty}_{R}(X)$. Then the sequence
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. Then there exists an $R_{G}$-projective resolution
[…]
with $\varepsilon(1_{X})=1_{X}$ and $\partial_{1}(1_{X}[i])=\alpha_{\varphi_{i}}-1_{\textup{im}\varphi_{i}}$. Here $1_{X}[i]\in\bigoplus_{i=1}^{n}R_{G}$ is the element such that the $i$-th coordinate is $1_{X}$ and the other coordinates are 00.
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. Then for every left $R_{G}$-module $M$, the cohomology group $\textup{H}^{0}(G,M)$ is naturally identified with the set
$\{\xi\in M\mid(\alpha_{\varphi_{i}}-1_{\textup{im}\varphi_{i}})\xi=0\ \forall i=1,...,n\}.$ (4.2)
@@
4.2 Cohomological dimension
Recall that the $R$-cohomological dimension of $G$ is denoted by $\textup{cd}_{R}(G)$ (Definiton 1.14). Note that $\textup{cd}_{R}(G)=\textup{pd}_{R_{G}}(l^{\infty}_{R}(X))$ holds by Proposition 2.14, and $\textup{cd}_{R}(G)\leq\textup{cd}_{\mathbb{Z}}(G)$ holds by Lemma 4.5.
First we compute the cohomological dimension of some examples.
@@ 4.2 Cohomological dimension
$G=\left\{\left(x,s^{\pm 1}x\right)\in X\times X\>\middle|\>x\in X,\ s\in S\right\}.$
@@
$0\to l^{\infty}_{R}(X)\otimes_{R}P_{n}\to\cdots\to l^{\infty}_{R}(X)\otimes_{R}P_{0}\to l^{\infty}_{R}(X)\to 0$ (4.3)
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. We will show that the sequence
[…]
@@
$\bigsqcup_{i=1}^{n}\left\{\left(y,\varphi_{i}^{-1}y\right)\mid y\in[x]_{G}\cap\textup{im}\varphi_{i}\right\}$ (4.4)
@@
$\left(y,\varphi_{i}^{-1}(y)\right)\mapsto a_{i}(x,y)$
for $1\leq i\leq n$, and let $c_{x}$ be the finitely supported function on $[x]_{G}$ defined by
@@
[…]
@@
[…]
@@
[…]
@@
$-\circ\sigma:\textup{Hom}_{R_{G}}\left(\bigoplus_{i=1}^{n}R_{G},R_{G}\right)\to\textup{Hom}_{R_{G}}(\textup{im}\partial_{1},R_{G}).$
@@
Take $\{\varphi_{i}\}_{i=0}^{\infty}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that for every $k\in\mathbb{Z}_{\geq 0}$, there exists $n\in\mathbb{Z}_{\geq 0}$ such that $G^{k}=\bigcup_{i=1}^{n}\textup{graph}\varphi_{i}$. Then we define $\{\psi_{i}\}_{i=0}^{\infty}$ as follows: Set $\psi_{0}=\varphi_{0}$. For $i\geq 1$, let $\psi_{i}$ be the restriction of $\varphi_{i}$ such that
@@
We claim that this left $R_{G}$-module is projective. Indeed, take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G^{\prime}\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that $\bigsqcup_{i=1}^{n}\textup{dom}\varphi_{i}=Y$ and $\bigcup_{i=1}^{n}\textup{im}\varphi_{i}\subset\gamma(X)$. Then we have

### v3

@@ 3.1. Construction of trees
A family $\mathcal{C}$ of non-empty proper subsets of $X$ is a treeset on $X$ if it is
@@
    nested, i.e., for all $C,D\in\mathcal{C}$, either
[…]
@@
    closed under complementation, i.e., if $C\in\mathcal{C}$, then $\overline{C}\in\mathcal{C}$ holds, and
@@
    finitely separating, i.e., for all $x,y\in X$, we have
    $|\{C\in\mathcal{C}\mid x\in C,\ y\notin C\}|<\infty.$
The structure tree associated to a treeset is introduced by [Dun79]. We will present a construction of it, which is based on [DD89, Chapter II] but described from the viewpoint of ultrafilters in the sense of [Rol98]. The proofs are given for the reader’s convenience.
@@
[…]
@@
Next, we show that $(V_{\mathcal{C}},T_{\mathcal{C}})$ is acyclic. Let $\{u_{i}\}_{i=0}^{n}$ be a cycle of $T_{\mathcal{C}}$ with $n\geq 2$ such that the path $\{u_{i}\}_{i=0}^{n-1}$ is simple. Let $u_{0}\setminus u_{1}=\{C\}$. Then by Lemma 3.7, for $1\leq i\leq n-1$, we have $u_{i+1}\setminus u_{i}\neq\{C\}$ since $(u_{i+1},u_{i})\neq(u_{0},u_{1})$. This implies that $C\notin u_{i}$ for $1\leq i\leq n$ by the induction on $i$, which contradicts that $u_{n}=u_{0}$. Hence this graph is a tree. ∎
@@
3.2. Borel cutsets of Borel graphs
@@
Take a $G$-path $\{x_{i}\}_{i=0}^{n}$ from $x$ to $y$. If $C\in\mathcal{C}$ satisfies $x\in C$ and $y\notin C$, then we have $x_{i}\in\partial_{\textup{iv}}^{G}C$ for some $0\leq i\leq n-1$. By the previous lemma, there exist only finitely many $C\in\mathcal{C}$ satisfying this property. Hence the lemma is proved. ∎
The following is inspired by the proof of [CPT+25, Proposition 3.3]:
@@
Take Borel subset $\mathcal{C}^{+}\subset\mathcal{C}$ so that $\left|\left\{C,\overline{C}\right\}\cap\mathcal{C}^{+}\right|=1$ for every $C\in\mathcal{C}$. By [KST99, Proposition 4.6], the Borel graph
$\bigsqcup_{\omega\in X/E_{G}}\left\{(C,D)\in\mathcal{C}^{+}_{\omega}\times\mathcal{C}^{+}_{\omega}\>\middle|\>C\textup{ and }D\textup{ are non-nested with each other}\right\}$
@@
3.3. Structure trees for Borel graphs
@@
for every $0\leq i\leq n-1$. It suffices to verify that $x_{0}\neq x_{2n}$. Note that $\rho(x_{2i})\neq\rho(x_{2i+1})=\rho(x_{2i+2})$ for every $0\leq i\leq n-1$. We prove that
(3.2) $d_{T_{\mathcal{C}}}(\rho(x_{0}),\rho(x_{2n}))=\sum_{i=0}^{n-1}d_{T_{\mathcal{C}}}(\rho(x_{2i}),\rho(x_{2i+2}))$
@@
4. Cohomology of graphs
Througout this section, let $R$ be a non-zero commutative ring and $(X,G)$ an abstract graph with uniformly bounded degrees. Many discussions here are analogous to the usual group cohomology (see e.g., [Coh72]).
4.1. Computaion of cohomology
@@
For $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$, set $\alpha_{\varphi}=1_{\textup{graph}\varphi}\in R_{G}$. Note that
@@
hold for $\varphi,\psi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ and $f\in l^{\infty}_{R}(X)$, where $\alpha_{\varphi*}f$ is the multiplication in the left $R_{G}$-module $l^{\infty}_{R}(X)$ as defined in Definition 1.10.
@@
If $G=\bigcup_{i=1}^{n}\textup{graph}\varphi_{i}$ with $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$, then $R_{G}$ is generated by the set $\{\alpha_{\varphi_{i}}\}_{i=1}^{n}\cup l^{\infty}_{R}(X)$ as a ring. Indeed, for every $k\in\mathbb{Z}_{\geq 0}$, there exists $\{\psi_{j}\}_{j=1}^{m}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ such that $G^{k}=\bigcup_{j=1}^{m}\textup{graph}\psi_{j}$ and each $\psi_{j}$ is obtained by composing elements of $\{\varphi_{i}\}_{i=1}^{n}$ at most $k$ times. Then for every $a\in R_{G}^{k}$, there exists $\{f_{j}\}_{j=1}^{m}\subset l^{\infty}_{R}(X)$ such that
@@
The algebra $R_{G}$ is projective as a left $l^{\infty}_{R}(X)$-module. Indeed, by Remark 4.2, we can take $\{\varphi_{i}\}_{i=1}^{\infty}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that for every $k\in\mathbb{Z}_{\geq 0}$, there exists $n\in\mathbb{Z}_{\geq 1}$ such that $G^{k}=\bigsqcup_{i=1}^{n}\textup{graph}\varphi_{i}$. Then we have $R_{G}=\bigoplus_{i=1}^{\infty}l^{\infty}_{R}(X)\cdot\alpha_{\varphi_{i}}$, where each $l^{\infty}_{R}(X)\cdot\alpha_{\varphi_{i}}$ is projective since we have
@@
(4.1) $\cdots\to l^{\infty}_{R}(X)\otimes_{l^{\infty}_{\mathbb{Z}}(X)}P_{n}\to\cdots\to l^{\infty}_{R}(X)\otimes_{l^{\infty}_{\mathbb{Z}}(X)}P_{0}\to l^{\infty}_{R}(X)\to 0$
@@
for $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ and $f\in l^{\infty}_{R}(X)$. Then the sequence
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. Then there exists an $R_{G}$-projective resolution
$\cdots\to\bigoplus_{i=1}^{n}R_{G}\xrightarrow{\partial_{1}}R_{G}\xrightarrow{\varepsilon}l^{\infty}_{R}(X)\to 0$
with $\varepsilon(1_{X})=1_{X}$ and $\partial_{1}(1_{X}[i])=\alpha_{\varphi_{i}}-1_{\textup{im}\varphi_{i}}$. Here $1_{X}[i]\in\bigoplus_{i=1}^{n}R_{G}$ is the element such that the $i$-th coordinate is $1_{X}$ and the other coordinates are 0.
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. Then for every left $R_{G}$-module $M$, the cohomology group $\textup{H}^{0}(G,M)$ is naturally identified with the set
(4.2) $\{\xi\in M\mid(\alpha_{\varphi_{i}}-1_{\textup{im}\varphi_{i}})\xi=0\ \forall i=1,...,n\}.$
@@
4.2. Cohomological dimension
Recall that the $R$-cohomological dimension of $G$ is denoted by $\textup{cd}_{R}(G)$ (Definiton 1.15). Note that $\textup{cd}_{R}(G)=\textup{pd}_{R_{G}}(l^{\infty}_{R}(X))$ holds by Proposition 2.14, and $\textup{cd}_{R}(G)\leq\textup{cd}_{\mathbb{Z}}(G)$ holds by Lemma 4.5. First we compute the cohomological dimension of some examples.
@@ 4.2. Cohomological dimension
$G=\left\{(x,s^{\pm 1}x)\in X\times X\>\middle|\>x\in X,\ s\in S\right\}.$
@@
(4.3) $0\to l^{\infty}_{R}(X)\otimes_{R}P_{n}\to\cdots\to l^{\infty}_{R}(X)\otimes_{R}P_{0}\to l^{\infty}_{R}(X)\to 0$
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. We will show that the sequence
$0\to\bigoplus_{i=1}^{n}R_{G}\cdot 1_{\textup{im}\varphi_{i}}\xrightarrow{\partial_{1}}R_{G}\xrightarrow{\varepsilon}l^{\infty}_{R}(X)\to 0$
@@
(4.4) $\bigsqcup_{i=1}^{n}\left\{(y,\varphi_{i}^{-1}y)\mid y\in[x]_{G}\cap\textup{im}\varphi_{i}\right\}$
@@
$(y,\varphi_{i}^{-1}(y))\mapsto a_{i}(x,y)$
for $1\leq i\leq n$, and let $c_{x}$ be the finitely supported function on $x]_{G}$ defined by
@@
To show the converse, suppose that $\textup{cd}_{R}(G)=0$. Then since the left $R_{G}$-module $l^{\infty}_{R}(X)$ is projective, we can take a semi-inverse of the map $\varepsilon:R_{G}\to l^{\infty}_{R}(X),\ 1_{X}\mapsto 1_{X}$, that is, there exists $\sigma\in\textup{Hom}_{R_{G}}(l^{\infty}_{R}(X),R_{G})$ such that $\varepsilon\circ\sigma=\textup{id}_{l^{\infty}_{R}(X)}$. Fix $x\in X$. Then for all $y,z\in[x]_{G}$, we have
@@
$\cdots\to\bigoplus_{i=1}^{n}R_{G}\xrightarrow{\partial_{1}}R_{G}\xrightarrow{\varepsilon}l^{\infty}_{R}(X)\to 0$
@@
$0\to\textup{im}\partial_{1}\hookrightarrow R_{G}\xrightarrow{\varepsilon}l^{\infty}_{R}(X)\to 0.$
@@
$\circ\sigma:\textup{Hom}_{R_{G}}\left(\bigoplus_{i=1}^{n}R_{G},R_{G}\right)\to\textup{Hom}_{R_{G}}(\textup{im}\partial_{1},R_{G}).$
@@
Take $\{\varphi_{i}\}_{i=0}^{\infty}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that for every $k\in\mathbb{Z}_{\geq 0}$, there exists $n\in\mathbb{Z}_{\geq 0}$ such that $G^{k}=\bigcup_{i=1}^{n}\textup{graph}\varphi_{i}$. Then we define $\{\psi_{i}\}_{i=0}^{\infty}$ as follows: Set $\psi_{0}=\varphi_{0}$. For $i\geq 1$, let $\psi_{i}$ be the restriction of $\varphi_{i}$ such that
@@
We claim that this left $R_{G}$-module is projective. Indeed, take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G^{\prime}\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that $\bigsqcup_{i=1}^{n}\textup{dom}\varphi_{i}=Y$ and $\bigcup_{i=1}^{n}\textup{im}\varphi_{i}\subset\gamma(X)$. Then we have